OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)] −13 −25 25 17

OpenStudy (anonymous):

first you have to find g(x) since it gave you what x is for g. In the equation that was given to you for g(x) where ever there is a x put a (-2) and figure it out and tell me what you get.

OpenStudy (anonymous):

17

OpenStudy (anonymous):

\[g(-2)=6(-2)-1\] \[g(-2)=-12-1\] \[g(-2)=-13\] now you put -13 into x for f(x) \[f(-13)=-2(-13)-1\] \[f(-13)=26-1\] \[f(-13)=25\]